# On the conditioning of histopolationL. Bruni Bruno, S. Serra-Capizzano

### v1

@@ 3.3 Conditioning with respect to the Frobenius norm
Proposition 3.1 and subsequent computations allow to deduce results beyond the conditioning in the Euclidean norm. Let us recall that, for a matrix $A\in\mathbb{R}^{m\times n}$, its Forbenius norm is the quantity
$\|A\|_{F}:=\sqrt{\sum_{i=1}^{m}\sum_{i=1}^{n}|A_{i,j}|^{2}}$
and the conditioning in the Frobenius norm is thus
(16) $\kappa_{F}(A):=\|A\|_{F}\|A^{-1}\|_{F}.$
The quantity $\kappa_{F}(A)$ reflects some geometrical features of the matrix $A$ [7], and in the context of nodal interpolation with Chebyshev nodes it has been studied in [9]. For large $d$, we are in a position to provide an asymptotic trend for $\kappa_{F}(V_{d})$. To this end, we shall need the integral sine function
@@
One has
(17) $\lim_{d\to\infty}\|V_{d}\|_{F}=\sqrt{2a\mathrm{Si}(2a)+\cos(2a)-1}.$
In particular, if $a=\pi/2$, one has $\lim_{d\to\infty}\|V_{d}\|_{F}\leq 2$.
Proof 3.11.
First of all, recall that $\|V_{d}\|_{F}^{2}=\mathrm{Tr}\left(V_{d}^{\top}V_{d}\right)$. Hence, by Proposition 3.1, we get
$\|V_{d}\|_{F}^{2}=\mathrm{Tr}\left(V_{d}^{\top}V_{d}\right)=\sum_{i=1}^{d}\frac{2d}{i^{2}}\sin^{2}(i\rho)=\sum_{i=1}^{d}\frac{2d}{i^{2}}\sin^{2}\left(i\frac{a}{d}\right).$
The terms of this sum are positive and decreasing to 0, and the sum converges due to the presence of the term $i^{2}$. By the definition of Riemann integral, setting $\Delta x:=1/d$ and $x_{i}:=i/d$ and passing to the limit for $d\to\infty$, one has
$\lim_{d\to+\infty}\sum_{i=1}^{d}\frac{2d}{i^{2}}\sin^{2}\left(a\frac{i}{d}\right)=\lim_{d\to+\infty}\sum_{i=1}^{d}\frac{1}{x_{i}^{2}}\sin^{2}\left(ax_{i}\right)\Delta x=2\int_{0}^{1}\frac{\sin^{2}\left(ax\right)}{x^{2}}\mathrm{d}x.$
This integral can be expressed using the sine integral function $\mathrm{Si}(z)$, and after the immediate change of variable $t=ax$ and applying integration by parts, one obtains
$2\int_{0}^{1}\frac{\sin^{2}\left(ax\right)}{x^{2}}\mathrm{d}x=2a\int_{0}^{a}\frac{\sin^{2}(t)}{t^{2}}\mathrm{d}t=2a\mathrm{Si}(2a)+\cos(2a)-1.$
Taking the square root we get the claim, and substituting $a=\pi/2$ we immediately obtain the respective bound, as $\mathrm{Si}(\pi)\approx 1.85$.
Lemma 3.10 gives an exact formula for the first term of the right hand side of (16). Althouhgh such a neat result cannot be replicated for $\|V_{d}^{-1}\|_{F}$, we can at least estimate its trend as $d$ grows large.
@@
Let $0<a\leq\pi/2$. One has
$\left\|V_{d}^{-1}\right\|_{F}\leq\frac{d}{\sqrt{2}a}.$
In particular, for $a=\pi/2$, one has $\|V_{d}^{-1}\|_{F}\leq d\frac{\sqrt{2}}{\pi}$.
Proof 3.13.
Again, let us observe that $\|V_{d}^{-1}\|^{2}_{F}=\mathrm{Tr}\left((V_{d}^{-1})^{\top}V_{d}^{-1}\right)=\mathrm{Tr}\left(V_{d}^{-\top}V_{d}^{-1}\right)$. Recalling formula (14), which states that $V_{d}^{-1}=W_{d}^{-1}V_{d}^{\top}$, being $W_{d}$ the diagonal matrix $W_{d}=V_{d}^{\top}V_{d}$, exploiting the cyclic property of the trace we get
$\left\|V_{d}^{-1}\right\|^{2}_{F}$ $=\mathrm{Tr}\left(V_{d}^{-\top}V_{d}^{-1}\right)=\mathrm{Tr}\left(V_{d}^{-\top}W_{d}^{-1}V_{d}^{\top}\right)=\mathrm{Tr}\left(V_{d}^{\top}V_{d}^{-\top}W_{d}^{-1}\right)=\mathrm{Tr}\left(W_{d}^{-1}\right).$
By Proposition 3.1, $W_{d}^{-1}$ is a diagonal matrix whose diagonal entries are the reciprocal of the diagonal entries of $V_{d}^{\top}V_{d}$, that is,
$\left[W_{d}^{-1}\right]_{i,i}=\frac{i^{2}}{2d\sin^{2}\left(\frac{a}{d}\right)}.$
Now, observe that, for $0\leq t\leq\pi/2$ one has $\sin(t)\geq t/2$. Since $a\leq\pi/2$ and $i/d\leq 1$, we may bound
$\mathrm{Tr}\left(W_{d}^{-1}\right)=\sum_{i=1}^{d}\frac{i^{2}}{2d\sin^{2}(a\frac{i}{d})}\leq\sum_{i=1}^{d}\frac{i^{2}}{2d(a\frac{i}{d})^{2}}=\sum_{i=1}^{d}\frac{d}{2a^{2}}=\frac{d^{2}}{2a^{2}}.$
By taking the square root of this quantity, we immediately get that $\|V_{d}^{-1}\|_{F}\leq\frac{d}{\sqrt{2}a}$.
We are in a position to prove the main result of the section.
@@
The conditioning $\kappa_{F}(V_{d})$ grows asymptotically linearly with $d$.
Proof 3.15.
Lemma 3.10 gives an equality for the term $\|V_{d}\|_{F}$, while Lemma 3.12 gives an upper bound for $\|V_{d}^{-1}\|_{F}$. Since $\kappa_{F}(V_{d})=\|V_{d}\|_{F}\|V_{d}^{-1}\|_{F}$, the linear upper bound is established.
To obtain the lower bound, it is sufficient to observe that
$\sum_{i=1}^{d}\frac{i^{2}}{2d\sin^{2}\left(\frac{i}{d}a\right)}\geq\frac{1}{2d}\sum_{i=1}^{d}i^{2}=\frac{1}{2}\frac{d(d+1)(2d+1)}{6}\geq\frac{2d^{3}}{12d}=\frac{d^{2}}{6}.$
The square root of this quantity offers thus a linear lower bound for $\|V_{d}^{-1}\|_{F}$, whence the claim.
The result of Theorem 3.14 is depicted, for the limiting case $a=\pi/2$, in Figure 9.
Figure 9: Linear conditioning in the Frobenius norm of the Vandermonde matrix in the Chebyshev basis for classes (C2).
@@
Although the conditioning with respect to the Euclidean norm is the most common in numerical analysis, it is important to point out that it is not the only one. For instance, the conditioning with respect to the $\infty$-norm is usually taken into account when dealing with Bernstein-Vandermonde collocation matrices, whose numerical features are not only of theoretical interest, but also have several applications [1]. Thus, a natural extension to this work consists in considering the Bernstein basis (fixing the reference interval to I=[1,1]I=[-1,1]) and analyzing the relative conditioning in the $\infty$-norm. A well-established and interesting starting point for a worthwhile comparison with the literature is represented by the work [10].

### v2

@@ 3.3 Conditioning with respect to the Frobenius norm
Let us recall that, for a matrix $A\in\mathbb{R}^{m\times n}$, its Forbenius norm is the quantity
$\|A\|_{F}:=\sqrt{\sum_{i=1}^{m}\sum_{i=1}^{n}|A_{i,j}|^{2}}=\sqrt{\mathrm{Tr}(A^{\top}A)}$
and the corresponding condition number is thus
$\kappa_{F}(A):=\|A\|_{F}\|A^{-1}\|_{F}.$ (28)
The quantity $\kappa_{F}(A)$ reflects some geometrical features of the matrix $A$ [9], and in the context of nodal interpolation with Chebyshev nodes it has been studied in [11]. The relationship between $\kappa_{2}(A)$ and $\kappa_{F}(A)$ is expressed by the inequality
$\kappa_{F}(A)\geq\kappa_{2}(A)+\frac{1}{\kappa_{2}(A)}+m-2,$
see [2], which implies that the ill-conditioning associated with the monomial basis and proved in Theorem 2.5 immediately carries over to the Frobenius setting.
3.3.1 The Frobenius conditioning for Chebyshev segments
Proposition 3.1 and subsequent computations allow to deduce results beyond the conditioning in the Euclidean norm for the histopolation matrix $V_{d}$. First, observe that the Cauchy-Schwarz inequality applied to (28) gives
$\kappa_{F}(V_{d})=\|V_{d}\|_{F}\|V_{d}^{-1}\|_{F}=\sqrt{\sum_{i=1}^{d}\sigma_{i}^{2}(V_{d})}\sqrt{\sum_{i=1}^{d}\frac{1}{\sigma_{i}^{2}(V_{d})}}\geq\sum_{i=1}^{d}1=d,$ (29)
which yields a linear lower bound for $\kappa_{F}(V_{d})$. For large $d$, we are also in a position to provide an asymptotic upper bound for $\kappa_{F}(V_{d})$. To this end, we shall need the integral sine function
@@ 3.3.1 The Frobenius conditioning for Chebyshev segments
Let $0<a<1$. One has
$\lim_{d\to\infty}\|V_{d}\|_{F}=\sqrt{2a\pi\mathrm{Si}(2a\pi)-2\sin^{2}(a\pi)}.$ (30)
Proof.
First of all, recall that $\|V_{d}\|_{F}^{2}=\mathrm{Tr}\left(V_{d}^{\top}V_{d}\right)$. Hence, by Proposition 3.1, substituting $\rho=a\pi/d$, we get
$\|V_{d}\|_{F}^{2}=\mathrm{Tr}\left(V_{d}^{\top}V_{d}\right)=\sum_{i=1}^{d-1}\frac{2d}{i^{2}}\sin^{2}(i\rho)+\frac{4}{d}\sin^{2}(d\rho)=\sum_{i=1}^{d}\frac{2d}{i^{2}}\sin^{2}\left(i\frac{a}{d}\pi\right)+\frac{2}{d}\sin^{2}(a\pi).$
The terms of this sum are positive and decreasing to 0, and the sum converges due to the presence of the term $i^{2}$. We may hence interpret it as a Riemann integral where $\Delta x:=1/d$ and $x_{i}:=i/d$ and passing to the limit for $d\to\infty$. Thus, one has
$\lim_{d\to+\infty}\left(\sum_{i=1}^{d}\frac{2d}{i^{2}}\sin^{2}\left(i\frac{a}{d}\pi\right)+\frac{2}{d}\sin^{2}(a\pi)\right)=\lim_{d\to+\infty}\sum_{i=1}^{d}\frac{2d}{i^{2}}\sin^{2}\left(a\frac{i}{d}\pi\right)$
$=$ $\lim_{d\to+\infty}\sum_{i=1}^{d}\frac{2d^{2}}{i^{2}d}\sin^{2}\left(a\frac{i}{d}\pi\right)=\lim_{d\to+\infty}2\sum_{i=1}^{d}\frac{1}{x_{i}^{2}}\sin^{2}\left(a\pi x_{i}\right)\Delta x=2\int_{0}^{1}\frac{\sin^{2}\left(a\pi x\right)}{x^{2}}\mathrm{d}x.$
This integral can be expressed using the sine integral function $\mathrm{Si}(z)$, and after the immediate change of variable $t=a\pi x$ and applying integration by parts, one obtains
$2\int_{0}^{1}\frac{\sin^{2}\left(a\pi x\right)}{x^{2}}\mathrm{d}x$ $=2a\pi\int_{0}^{a\pi}\frac{\sin^{2}(t)}{t^{2}}\mathrm{d}t=2a\pi\left(\mathrm{Si}(2a\pi)-\frac{\sin^{2}(a\pi)}{a\pi}\right)$
$=2a\pi\mathrm{Si}(2a\pi)-2\sin^{2}(a\pi).$
Taking the square root we get the claim.∎
The quantity (30) can be numerically estimated, thus Lemma 3.18 can concretely be used to describe the asymptotics of the first term of the right hand side of (28). Further, due to the monotonicity as a function of $a$ for $0<a<1$, one concludes that
$\lim_{d\to\infty}\|V_{d}\|_{F}\leq\sqrt{2\pi\mathrm{Si}(\pi)},$
obtaining an upper bound for $\|V_{d}\|_{F}$ which is independent of $a$.
To obtain an upper bound for $\|V_{d}^{-1}\|_{F}$ and hence an estimate for the remaining term of (28), we exploit the explicit expression for $V_{d}^{-1}$ derived in Section 3.1.
@@
Let $0<\rho<\pi/d$. One has
$\left\|V_{d}^{-1}\right\|_{F}=\sqrt{\frac{1}{2d}\sum_{i=1}^{d-1}\left(\frac{i}{\sin(i\rho)}\right)^{2}+\frac{d}{4\sin^{2}(d\rho)}},$
whence, for $0<\rho<\pi/d$,
$\left\|V_{d}^{-1}\right\|_{F}\leq\frac{d}{\sqrt{2}\sin(d\rho)}=\frac{d}{\sqrt{2}\sin(a\pi)}.$ (31)
Proof.
By applying the definition of Frobenius norm, and working with squared quantities,
$\left\|V_{d}^{-1}\right\|_{F}^{2}=\sum_{i=1}^{d}\sum_{j=1}^{d}\left[V_{i,j}^{-1}\right]^{2}=\sum_{i=1}^{d-1}\sum_{j=1}^{d}\left[V_{i,j}^{-1}\right]^{2}+\sum_{j=1}^{d}\left[V_{d,j}^{-1}\right]^{2}$
whence, plugging the explicit terms given in Section 3.1, one finds
$\left\|V_{d}^{-1}\right\|_{F}^{2}$ $=\sum_{i=1}^{d-1}\sum_{j=1}^{d}\left(\frac{i\sin(i\tau_{j})}{d\sin(i\rho)}\right)^{2}+\sum_{j=1}^{d}\left(\frac{\sin(d\tau_{j})}{2\sin(d\rho)}\right)^{2}$
$=\sum_{i=1}^{d-1}\frac{i^{2}}{d^{2}\sin^{2}(i\rho)}\sum_{j=1}^{d}\sin^{2}(i\tau_{j})+\frac{1}{4\sin^{2}(d\rho)}\sum_{j=1}^{d}\sin^{2}(d\tau_{j}).$
Both sums depending on $j$ have been computed in (11). In particular, the first equals $d/2$, while the second equals $d$. Substistuting, this gives
$\left\|V_{d}^{-1}\right\|_{F}^{2}=\sum_{i=1}^{d-1}\frac{i^{2}}{d^{2}\sin^{2}(i\rho)}\frac{d}{2}+\frac{1}{4\sin^{2}(d\rho)}d=\frac{1}{2d}\sum_{i=1}^{d-1}\frac{i^{2}}{\sin^{2}(i\rho)}+\frac{d}{4\sin^{2}(d\rho)}.$
Taking the square root one has the first equality. Exploiting the concavity of the sine function in $(0,\pi)$, Lemma 3.6 extends to $j=d$ and we get that $\sin(x)/x\geq\sin(d\rho)/(d\rho)$ in $(0,d\rho)$, whence
$\frac{i}{\sin(i\rho)}\leq\frac{d}{\sin(d\rho)}$
for all the terms in the remaining sum. We may thus expand the above equation to get
$\left\|V_{d}^{-1}\right\|_{F}^{2}$ $=\frac{1}{2d}\sum_{i=1}^{d-1}\frac{i^{2}}{\sin^{2}(i\rho)}+\frac{d}{4\sin^{2}(d\rho)}\leq\frac{1}{2d}\sum_{i=1}^{d-1}\frac{d^{2}}{\sin^{2}(d\rho)}+\frac{d}{4\sin^{2}(d\rho)}$
$=\frac{d(d-1)}{2\sin^{2}(d\rho)}+\frac{d}{4\sin^{2}(d\rho)}=\frac{2d^{2}-d}{4\sin^{2}(d\rho)}\leq\frac{d^{2}}{2\sin^{2}(d\rho)}.$
Taking the square root and substituting $\rho=a\pi/d$, one also obtains $\left\|V_{d}^{-1}\right\|_{F}\leq\frac{d}{\sqrt{2}\sin(a\pi)}$. ∎
Merging Lemma 3.18 and Lemma 3.19 with the general lower bound (29), we readily obtain the main result of the section.
@@
The conditioning $\kappa_{F}(V_{d})$ grows asymptotically linearly with $d$. In particular, for $0<a<1$, one has
$d\leq\kappa_{F}(V_{d})\leq\frac{\sqrt{a\pi\mathrm{Si}(2a\pi)-\sin^{2}(a\pi)}}{\sin(a\pi)}d.$ (32)
Figure 11 depicts the trend of the Frobenius conditioning, predicted by multiplying the bounds given in Lemma 3.18 and Lemma 3.19, which appears indeed linear. For small $a$, by performing a Taylor expansion, one sees that the right hand side of (32) is indeed close to $d$.
Figure 11: Linear conditioning in the Frobenius norm of the Vandermonde matrix in the Chebyshev basis for $a=0.1$ and $a=0.5$, with the respective upper bounds.
@@
Although the conditioning with respect to the Euclidean norm is the most common in numerical analysis, it is important to point out that it is not the only one. For instance, the conditioning with respect to the $\infty$-norm is usually taken into account when dealing with Bernstein-Vandermonde collocation matrices [10], whose numerical features are not only of theoretical interest, but also have several applications [1]. Thus, a natural extension to this work consists in considering the Bernstein basis (fixing the reference interval to I=[1,1]I=[-1,1]) and analyzing the relative conditioning in the $\infty$-norm. A well-established and interesting starting point for a worthwhile comparison with the literature is represented by the work [12].